Check continuity


f(x) = e^1/x - 1/(e^1/x +1)  , x is nt equal to 0


         0  , x=0


 

Grade:12

1 Answers

Askiitians Expert Soumyajit IIT-Kharagpur
13 years ago

Dear Shefali Sharma,

Ans:- f(x)=e^1/x - 1/(e^1/x+1)

Lt x tends to 0 from + i.e x=0+h=∞

Lt x tends to 0 from negative i.e x=0-h= -1

Hence it is dicotinuous at x=0 

Please feel free to post as many doubts on our discussion forum as you can. If you find any question
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We
are all IITians and here to help you in your IIT JEE preparation.


All the best Shefali Sharma !!!

 


Regards,

Askiitians Experts
Soumyajit Das IIT Kharagpur